# Extended Circular Nim

### v1

@@ 2.4 Extended circular nim with seven piles
Proof.
Next, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\in P$. From the definition of $P$, for any position $M\in P$, $5\times\max(\{n_{i}\})=\sum{n_{i}}$ holds.
Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6})$ be an option of $M$. If $n^{\prime}_{1}=n_{1},n^{\prime}_{2}=n_{2},n^{\prime}_{5}=n_{5},$ or $n^{\prime}_{6}=n_{6},$ then $\max(\{n_{i}\})=\max(\{n^{\prime}_{i}\})$, but $\sum{n^{\prime}_{i}}<\sum{n_{i}}$. Therefore, $5\times\max(\{n^{\prime}_{i}\})>\sum{n^{\prime}_{i}}$ and $M^{\prime}\not\in P$.
If $n^{\prime}_{1}<n_{1},n^{\prime}_{2}<n_{2},n^{\prime}_{5}<n_{5},$ and $n^{\prime}_{6}<n_{6},$ then $n^{\prime}_{3}=n_{3}$ and $n^{\prime}_{4}=n_{4}$.
Since $n^{\prime}_{3}+n^{\prime}_{4}=n_{3}+n_{4}=n_{1}>n^{\prime}_{1}$, we do not have $n^{\prime}_{0}=0,n^{\prime}_{1}=n^{\prime}_{2}=n^{\prime}_{3}+n^{\prime}_{4}=n^{\prime}_{5}=n^{\prime}_{6}$.
Also, we do not have $n^{\prime}_{1}=0,n^{\prime}_{2}=n^{\prime}_{3}=n^{\prime}_{4}+n^{\prime}_{5}=n^{\prime}_{6}=n^{\prime}_{0}$ or $n^{\prime}_{2}=0,n^{\prime}_{3}=n^{\prime}_{4}=n^{\prime}_{5}+n^{\prime}_{6}=n^{\prime}_{0}=n^{\prime}_{1}$ or $n^{\prime}_{5}=0,n^{\prime}_{6}=n^{\prime}_{0}=n^{\prime}_{1}+n^{\prime}_{2}=n^{\prime}_{3}=n^{\prime}_{4}$ or $n^{\prime}_{6}=0,n^{\prime}_{0}=n^{\prime}_{1}=n^{\prime}_{2}+n^{\prime}_{3}=n^{\prime}_{4}=n^{\prime}_{5}$ because they contradict $n^{\prime}_{0}=n_{0}=0$ and $n^{\prime}_{3}+n^{\prime}_{4}>0$.
The case $n^{\prime}_{3}=0,n^{\prime}_{4}=n^{\prime}_{5}=n^{\prime}_{6}+n^{\prime}_{0}=n^{\prime}_{1}=n^{\prime}_{2}$ also contradicts because $n^{\prime}_{3}+n^{\prime}_{4}=n_{3}+n_{4}=n_{5}$ and $n^{\prime}_{4}=n_{5}>n^{\prime}_{5}$. In a similar way, $n^{\prime}_{4}=0,n^{\prime}_{5}=n^{\prime}_{6}=n^{\prime}_{0}+n^{\prime}_{1}=n^{\prime}_{2}=n^{\prime}_{3}$ contradicts, too.
Therefore, for $M\in P,$ every option $M^{\prime}\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\not\in N$.
Without loss of generality, we assume that $\min(\{n_{i}\})=n_{0}$. We consider three cases:
  • (i)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{1}$. $M^{\prime}=(n_{0},n_{1},n_{1},0,n_{1},n_{1},n_{1}-n_{0})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{6}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
  • (ii)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{2}$. $M^{\prime}=(n_{0},n_{2}-n_{0},n_{2},n_{2},0,n_{2},n_{2})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{5}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
  • (iii)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{3}$.
    • (iii-1)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{i}(i\in\{1,2,5\})$, $M^{\prime}=(0,n_{i},n_{i},n_{3},n_{i}-n_{3},n_{i},n_{i})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-2)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{6}$ and $n_{0}+n_{6}\leq\min(\{n_{1},n_{2},n_{4},n_{5}\})$, $M^{\prime}=(n_{0},n_{0}+n_{6},n_{0}+n_{6},0,n_{0}+n_{6},n_{0}+n_{6},n_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-3)
      Consider the case that $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{6}$ and $n_{0}+n_{6}>\min(\{n_{1},n_{2},n_{4},n_{5}\})$. Let $a=\min(\{n_{1},n_{2},n_{4},n_{5}\})$. $n_{0}\leq a,n_{6}\leq a,n_{0}+n_{6}>a$ hold.
      • (iii-3-a)
        If $n_{1}=a$, $n_{2}=a$, or $n_{5}=a$, $M^{\prime}=(n_{0},a,a,0,a,a,a-n_{0})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (iii-3-b)
        If $n_{4}=a$, $M^{\prime}=(a-n_{6},a,a,0,a,a,n_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-4)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{4}$ and $n_{3}+n_{4}\leq\min(\{n_{1},n_{2},n_{5},n_{6}\})$, $M^{\prime}=(0,n_{3}+n_{4},n_{3}+n_{4},n_{3},n_{4},n_{3}+n_{4},n_{3}+n_{4})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-5)
      Consider the case that $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{4}$ and $n_{3}+n_{4}>\min(\{n_{1},n_{2},n_{5},n_{6}\})$. Let $a=\min(\{n_{1},n_{2},n_{5},n_{6}\})$. $n_{3}\leq a,n_{4}\leq a,n_{3}+n_{4}>a$ hold.
      • (iii-5-a)
        If $n_{2}=a$, $n_{5}=a$, or $n_{6}=a$, then $M^{\prime}=(0,a,a,a-n_{4},n_{4},a,a)$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (iii-5-b)
        If $n_{1}=a$, $M^{\prime}=(0,a,a,n_{3},a-n_{3},a,a)$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{4}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
@@
2.5 Extended circular nim with eight piles
@@
When $S=\{2\},$ the position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7},n_{8})$ can be considered as a disjunctive sum of positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in circular nim ${\rm CN}(4,i)$. If $i=2$, $M$ is disjunctive sum of two positions in ${\mathrm{C}N}(4,2)$, but any closed formula for Sprague-Grundy values of ${\mathrm{C}N}(4,2)$ is unknown. Therefore, it is an open problem to characterize $\mathcal{P}$-positions in ${\rm ECN}(8_{\{2\}},2)$. If $i=3,$ $M$ is disjunctive sum of two positions in ${\mathrm{C}N}(4,3)$. The formula for Sprague-Grundy value is shown in [7], thus this case is solved. If $i\geq 4,$ $M$ is the disjunctive sum of positions in ${\mathrm{C}N}(4,i)$. In this case, each player can choose all four piles which have $n_{0},n_{2},n_{4},n_{6}$ tokens or all four piles which have $n_{1},n_{3},n_{5},n_{7}$ tokens. Therefore, this position can be considered as a position $(n_{0}+n_{2}+n_{4}+n_{6},n_{1}+n_{3}+n_{5}+n_{7})$ in two-pile nim and $M$ is a $\mathcal{P}$-position if and only if $n_{0}+n_{2}+n_{4}+n_{6}=n_{1}+n_{3}+n_{5}+n_{7}$.
When $S=\{3\}$ ${\rm ECN}(8_{3},i)$ is isomorphic to ${\rm ECN}(8_{1},i)$. As shown in Figure 4, we have this isomorphism by exchanging names of piles.
@@
When $S=\{4\}$ and $i\geq 2$, ${\rm ECN}(8_{\{4\}},i)$ can be considered as a position $(n_{0}+n_{4},n_{1}+n_{5},n_{2}+n_{6},n_{3}+n_{7})$ in four-pile nim and $M$ is a $\mathcal{P}$-position if and only if $(n_{0}+n_{4})\oplus(n_{1}+n_{5})\oplus(n_{2}+n_{6})\oplus(n_{3}+n_{7})=0$
Theorem 13.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,3\}},2)$. $M$ is a $\mathcal{P}$-position if and only if $M\in S=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}\oplus n_{2}\oplus n_{4}\oplus n_{6}=n_{1}\oplus n_{3}\oplus n_{5}\oplus n_{7}=0\}$.
Proof.
In ${\rm ECN}(8_{\{1,3\}},2)$, the player can reduce at most one of $n_{0},n_{2},n_{4},n_{6}$, and at most one of $n_{1},n_{3},n_{5},n_{7}$ in one move. Therefore, this ruleset is selective sum of two positions $(n_{0},n_{2},n_{4},n_{6}),$ and $(n_{1},n_{3},n_{5},n_{7})$ in four-pile nim. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such two four-pile nim positions are $\mathcal{P}$-positions. ∎
@@
In ${\rm ECN}(8_{\{1,3\}},4)$, the player can reduce at most one of $n_{0},n_{4}$, at most one of $n_{1},n_{5}$, at most one of $n_{2},n_{6},$ and at most one of $n_{3},n_{7}$ in one move. Therefore, this ruleset is the selective sum of four positions $(n_{0},n_{4}),(n_{1},n_{5}),(n_{2},n_{6}),$ and $(n_{3},n_{7})$ in two-pile nim. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such four two-pile nim positions are $\mathcal{P}$-positions. ∎
@@
Proof.
In ${\rm ECN}(8_{\{1,3\}},6)$, the player can reduce at most three of $n_{0},n_{2},n_{4},n_{6}$ and at most three of $n_{1},n_{3},n_{5},n_{7}$ in one move. Therefore, this ruleset is the selective sum of two positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in ${\rm MN}(4,3)$. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such two Moore’s nim positions are $\mathcal{P}$-positions. ∎
When $S=\{1,4\}$ and $i\leq 4$, ${\rm ECN}(8_{\{1,4\}},i)$ has not been solved. For the case of $i\geq 5$, ${\rm ECN}(8_{\{1,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1\}},i)$.
@@
When $S=\{2,4\}$, similar to the case of $S=\{2\}$, the position is a disjunctive sum of two positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in ${\rm ECN}(4_{\{1,2\}},\min(i,4))$.
@@
Consider a non-terminal position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\in P$.
@@
    Consider the case that the player selects $n_{0},n_{2},n_{4},n_{6}$ or part of them to reduce. Let $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5},n^{\prime}_{6},n_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$. If $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$, then $n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{0}$. Also, since $M\in P$ and $n_{0}=n_{2}=n_{4}=n_{6}$, we do not have $n_{1}=n_{3}=n_{5}=n_{7}$. Therefore, $M^{\prime}\in N$.
@@
    Consider the case that the player selects $n_{1},n_{3},n_{5},n_{7}$ or part of them to reduce. Let $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n^{\prime}_{3},n_{4},n^{\prime}_{5},n_{6},n^{\prime}_{7})(n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$. Since $n_{0}=n_{2}=n_{4}=n_{6}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, $M\in N$.
@@
    Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6},n^{\prime}_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6},n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$ after such a move. Then, we do not have $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Also, since $\max(n^{\prime}_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})=n_{0}=n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, we do not have $n^{\prime}_{1}=n^{\prime}_{3}=n^{\prime}_{5}=n^{\prime}_{7}=n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}$. Therefore, $M\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
@@
If $n_{0}=n_{2}=n_{4}=n_{6}=n_{1}+n_{3}+n_{5}+n_{7},n_{1}=n_{3}=n_{5}=n_{7}$, then $M^{\prime}=(n_{1},n_{1},0,n_{3},0,n_{5},0,n_{7})$ is an option of $M$ and $M^{\prime}\in P$.
@@
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}\leq\min(\{n_{1},n_{3},n_{5},n_{7}\})$. If we have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0},n_{2},0,n_{4},0,n_{6},0)$ is an option of $M$ and $M^{\prime}\in P$. If we do not have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0}+n_{2}+n_{4}+n_{6},n_{2},n_{0}+n_{2}+n_{4}+n_{6},n_{4},n_{0}+n_{2}+n_{4}+n_{6},n_{6},n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$ and $M^{\prime}\in P$.
@@
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}>\min(\{n_{1},n_{3},n_{5},n_{7}\})$. Without loss of generality, $n_{1}=\min(\{n_{1},n_{3},n_{5},n_{7}\})$. For the case of $n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ since $n_{1}>n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\})$, we have $(n_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})$ such that $n_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}=n_{1},n^{\prime}_{2}\leq n_{2},n^{\prime}_{4}\leq n_{4},n^{\prime}_{6}\leq n_{6},$ and not satisfy $n_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Therefore, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{1},n^{\prime}_{4},n_{1},n^{\prime}_{6},n_{1})$ is an option of $M$ and $M^{\prime}\in P$.
    Similarly, for the cases of $n_{2}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),n_{4}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ and $n_{6}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ we have an option $M^{\prime}\in P$.
@@
2.6 Further generalization
From the results on ${\rm ECN}(5_{1},3)$ and ${\rm ECN}(7_{1,2},5)$, we have a generalization as follows:
Theorem 17.
Let $m(>1)$ be an integer and assume that $12m+$ 1 is a prime number. Consider a position $M=(n_{0},n_{1},\ldots,n_{2m})$ in ${\rm ECN}((2m+1)_{\{1,2,\ldots,m-1\}},2m-1)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{M\mid n_{0}=0,n_{1}=n_{2}=\cdots=n_{m-1}=n_{m}+n_{m+1}=n_{m+2}=\cdots=n_{2m}\}$.
Proof.
The proof is similar to the proof of Theorem 12. ∎
Also, we have a generalization from the results on ${\rm ECN}(4_{\{1\}},2),{\rm ECN}(6_{\{1,3\}},2),$ and ${\rm ECN}(8_{\{1,3\}},2)$, and a generalization from the results on ${\rm ECN}(4_{\{1\}},2)$ and ${\rm ECN}(8_{\{1,3\}},6)$.
Theorem 18.
Proof.
The proof is similar to the proof of Theorem 13. ∎
Theorem 19.
Proof.
The proof is similar to the proof of Theorem 15. ∎

### v2

@@ 4 Extended circular nim with seven piles
Note that Theorem 14 is also a generalization of the result on ${\rm ECN}(5_{\{1\}},3)$.
@@
5 Extended circular nim with eight piles
@@
When $S=\{2\},$ the position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7},n_{8})$ can be considered as a disjunctive sum of positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in circular nim ${\rm CN}(4,i)$. If $i=2$, $M$ is a disjunctive sum of two positions in ${\rm CN}(4,2)$, but any closed formula for Sprague-Grundy values of ${\rm CN}(4,2)$ is unknown. Therefore, it is an open problem to characterize $\mathcal{P}$-positions in ${\rm ECN}(8_{\{2\}},2)$. If $i=3,$ $M$ is a disjunctive sum of two positions in ${\rm MN}(4,3)$. Therefore, by using Theorem 5, the set of $\mathcal{P}$-positions can be characterized. If $i\geq 4,$ $M$ is the disjunctive sum of positions in ${\rm CN}(4,i)$. In this case, each player can choose all four piles which have $n_{0},n_{2},n_{4},n_{6}$ tokens or all four piles which have $n_{1},n_{3},n_{5},n_{7}$ tokens. Therefore, this position can be considered as a position $(n_{0}+n_{2}+n_{4}+n_{6},n_{1}+n_{3}+n_{5}+n_{7})$ in two-pile nim and $M$ is a $\mathcal{P}$-position if and only if $n_{0}+n_{2}+n_{4}+n_{6}=n_{1}+n_{3}+n_{5}+n_{7}$.
When $S=\{3\},$ ${\rm ECN}(8_{\{3\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1\}},i)$. As shown in Figure 7, we have this isomorphism by exchanging names of piles.
@@
When $S=\{4\}$ and $i\geq 2$, ${\rm ECN}(8_{\{4\}},i)$ can be considered as a position $(n_{0}+n_{4},n_{1}+n_{5},n_{2}+n_{6},n_{3}+n_{7})$ in four-pile nim and $M$ is a $\mathcal{P}$-position if and only if $(n_{0}+n_{4})\oplus(n_{1}+n_{5})\oplus(n_{2}+n_{6})\oplus(n_{3}+n_{7})=0$ .
When $S=\{1,2\},{\rm ECN}(8_{\{1,2\}},i)$ is unsolved for $2\leq i\leq 6.$
For ${\rm ECN}(8_{\{1,3\}},2),$ we can obtain the following corollary from Theorem 11.
Corollary 3.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,3\}},2)$. $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}\oplus n_{2}\oplus n_{4}\oplus n_{6}=n_{1}\oplus n_{3}\oplus n_{5}\oplus n_{7}=0\}$.
@@
In ${\rm ECN}(8_{\{1,3\}},4)$, the player can reduce at most one of $n_{0},n_{4}$, at most one of $n_{1},n_{5}$, at most one of $n_{2},n_{6},$ and at most one of $n_{3},n_{7}$ in one move. Therefore, this ruleset is the selective sum of four positions $(n_{0},n_{4}),(n_{1},n_{5}),(n_{2},n_{6}),$ and $(n_{3},n_{7})$ in two-pile nim. Thus, from Theorems 1 and 6, $M$ is a $\mathcal{P}$-position if and only if $n_{0}\oplus n_{4}=n_{1}\oplus n_{5}=n_{2}\oplus n_{6}=n_{3}\oplus n_{7}=0,$ that is, $n_{0}=n_{4},n_{1}=n_{5},n_{2}=n_{6},n_{3}=n_{7}.$
For ${\rm ECN}(8_{\{1,3\}},6),$ we show the following more generalized theorem.
@@
Proof.
Note that every odd number is prime to $2^{m}$. In ${\rm ECN}((2^{m})_{\{1,3,\ldots,h\}},2^{m}-2)$, the player can reduce at most $2^{m-1}-1$ numbers of $n_{0},n_{2},\ldots,n_{2^{m}-2}$ and at most $2^{m-1}-1$ numbers of $n_{1},n_{3},\ldots,n_{2^{m}-1}$ in one move. Therefore, this ruleset is the selective sum of two positions $(n_{0},n_{2},\ldots,n_{2^{m}-2})$ and $(n_{1},n_{3},\ldots,n_{2^{m}-1})$ in ${\rm MN}(2^{m-1},2^{m-1}-1)$. Thus, from Theorems 2 and 6, $M$ is a $\mathcal{P}$-position if and only if $n_{0}=n_{2}=\cdots=n_{2^{m}-2}$ and $n_{1}=n_{3}=\cdots=n_{2^{m}-1}$. ∎
Corollary 4.
@@
Note that Theorem 16 is also a generalization of the result on ${\rm ECN}(4_{\{1\}},2)$.
When $S=\{1,4\}$ and $2\leq i\leq 4$, ${\rm ECN}(8_{\{1,4\}},i)$ has not been solved. For the case of $i\geq 5$, ${\rm ECN}(8_{\{1,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1\}},i)$.
@@
When $S=\{2,4\}$, similar to the case of $S=\{2\}$, the position is a disjunctive sum of two positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in ${\rm ECN}(4_{\{1,2\}},\min(\{4,i\}))$. If $i=2$, then ${\rm ECN}(4_{\{1,2\}},2)$ is isomorphic to ${\rm MN}(4,2)$, but closed formula for Sprague-Grundy values of this ruleset is not known. Therefore, it is an open problem to characterize $\mathcal{P}$-positions in ${\rm ECN}(8_{\{2,4\}},2).$ If $i\geq 3$, ${\rm ECN}(8_{\{2,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{2\}},i)$.
@@
Consider a non-terminal position $M\in_{\circlearrowleft}P$. Without loss of generality, we assume that $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\in P$.
@@
    Consider the case that the player selects $n_{0},n_{2},n_{4},n_{6}$ or part of them to reduce.
    Let $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5},n^{\prime}_{6},n_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$. If $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$, then $n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{0}$. Also, since $M\in P$ and $n_{0}=n_{2}=n_{4}=n_{6}$, we do not have $n_{1}=n_{3}=n_{5}=n_{7}$. Therefore, $M^{\prime}\in N$.
@@
    Consider the case that the player selects $n_{1},n_{3},n_{5},n_{7}$ or part of them to reduce.
    Let $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n^{\prime}_{3},n_{4},n^{\prime}_{5},n_{6},n^{\prime}_{7})(n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$. Since $n_{0}=n_{2}=n_{4}=n_{6}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, $M^{\prime}\in N$.
@@
    Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6},n^{\prime}_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6},n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$ after such a move. Then, we do not have $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Also, since $\max(n^{\prime}_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})=n_{0}=n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, we do not have $n^{\prime}_{1}=n^{\prime}_{3}=n^{\prime}_{5}=n^{\prime}_{7}=n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}$. Therefore, $M^{\prime}\in N$.
@@
If $n_{0}=n_{2}=n_{4}=n_{6}=n_{1}+n_{3}+n_{5}+n_{7}$ and $n_{1}=n_{3}=n_{5}=n_{7}$, then $M^{\prime}=(n_{1},n_{1},0,n_{3},0,n_{5},0,n_{7})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
@@
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}\leq\min(\{n_{1},n_{3},n_{5},n_{7}\})$. If we have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0},n_{2},0,n_{4},0,n_{6},0)$ is an option of $M$ and $M^{\prime}\in P$. If we do not have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0}+n_{2}+n_{4}+n_{6},n_{2},n_{0}+n_{2}+n_{4}+n_{6},n_{4},n_{0}+n_{2}+n_{4}+n_{6},n_{6},n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
@@
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}>\min(\{n_{1},n_{3},n_{5},n_{7}\})$. Without loss of generality, $n_{1}=\min(\{n_{1},n_{3},n_{5},n_{7}\})$. For the case of $n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ since $n_{1}>n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\})$, we have $(n_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})$ such that $n_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}=n_{1},n^{\prime}_{2}\leq n_{2},n^{\prime}_{4}\leq n_{4},n^{\prime}_{6}\leq n_{6},$ and not satisfy $n_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Therefore, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{1},n^{\prime}_{4},n_{1},n^{\prime}_{6},n_{1})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    Similarly, for the cases of $n_{2}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),n_{4}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ and $n_{6}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ we have an option $M^{\prime}\in_{\circlearrowleft}P$.